OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

hartnn (hartnn):

let the age of Tony's age be x since combined age is 56 Belinda's age = 56-x Belinda is 8 more than twice Tony’s age. so 2*x+8=56-x can u solve this to get x ??

OpenStudy (anonymous):

tony's age=x blinda]s age=y x+y=56 y=2x+8 so combine the two previous equation and find the ages

OpenStudy (anonymous):

how would i combine them?

hartnn (hartnn):

like i did and u will get 2*x+8=56-x can u solve this to get x ??

OpenStudy (anonymous):

is it 2 times x? is that why you put the *

OpenStudy (anonymous):

yes

hartnn (hartnn):

yes,two times x *=multiplication

hartnn (hartnn):

let me give u hint to solve 2x+8=56-x 1. Add x to both sides 2. Subtract 8 from both sides what u get?

OpenStudy (anonymous):

would i subtract 8 from 8 and 56?

hartnn (hartnn):

yes,both

hartnn (hartnn):

so did u get something like this? 3x=48 ?

OpenStudy (anonymous):

x = 16

OpenStudy (anonymous):

?

hartnn (hartnn):

thats correct :) good work. so tony age = 16 Belinda age= 56-16 = ??

OpenStudy (anonymous):

40

hartnn (hartnn):

awesome :) u got both's age .

OpenStudy (anonymous):

how would i write the answer and thank you

hartnn (hartnn):

The age of Tony is 16 years, and the age of Belinda is 40 years.